OpenStudy (anonymous):

What is the product of 3the square root of u. (the square root of 448. + 4the square root of 7.)? Simplify if possible.

OpenStudy (jiteshmeghwal9):

Do u want to solve \[\sqrt{448}+4\sqrt{7}\]

OpenStudy (jiteshmeghwal9):

@erinjean ????

OpenStudy (anonymous):

3 square root of u. (the square root of 448. + 4the square root of 7.)

OpenStudy (jiteshmeghwal9):

\[3\times \sqrt{448}+4\times \sqrt{7}\]\[3\times8\sqrt{7}+4\sqrt{7}\]\[24\sqrt{7}+4\sqrt{7}=28\sqrt{7}\]